Hello, I'm having some trouble with this question with age, it is: Justine is 15years older than Emily now, but in 6 years time she will be twice as old as Emily. Write an equation and solve it to find the girls' present ages. I gave it ago using a method we learnt at school but it didn't quite work. Please help and thank you!
Answer by Alan3354(69443)

Justine is 15years older than Emily now, but in 6 years time she will be twice as old as Emily.
----------------
The age difference will always be 15 years.
--> 15 & 30 are 2x
Subtract 6
==============
J + 6 = 2(E + 6)
J = E + 15
---
E + 15 + 6 = 2(E + 6) = 2E + 12
--> E = 9, J = 24